OpenStudy (anonymous):

Use all nine operations and numbers once to make a calculation of 17. (,),+,-,x,2,4,5,7

OpenStudy (anonymous):

Please help me

OpenStudy (anonymous):

(2+0) = 2 + 1 = 3 - 2 = 1 x 17 = 17 + 2 = 19 + 7 = 24 - 5 = 19 - 2 = 17.

OpenStudy (anonymous):

Your supposed to use them all once.

OpenStudy (anonymous):

Then I'm not sure.

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

(7-4)x5+2=17

OpenStudy (anonymous):

Thanks anyways

OpenStudy (anonymous):

(7-4) = 3 x 5 = 15 + 2 = 17. Yep! That's correct!